OpenStudy (anonymous):

write a polynonmial function in standard form with the given zero's x= -2,0,1

OpenStudy (anonymous):

hint: If a function has 0 at x=a, x=b and x=c, it can be writen as f(x)=(x-a)(x-b)(x-c) now multiply the terms in paréntesis and you get the standart form

OpenStudy (anonymous):

f(x)=x-(-2)(x-0)(x-1)

OpenStudy (anonymous):

lie that

OpenStudy (anonymous):

f(x)=(x+2)(x-0)(x-1)

OpenStudy (anonymous):

now, multiply the parentesys

OpenStudy (anonymous):

x^2+2x

OpenStudy (anonymous):

hmm, let's do it step by step. f(x)=(x+2)(x-0)(x-1)=x(x+2)(x-1)=(x^2+2x)(x-1)=x^3-x^2+2x^2-2x=x^3+x^2-2x

OpenStudy (anonymous):

is the bottom one the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanx :).....do u think u could help me on a a couple moree